Calculate the molarity of a sucrose solution made by dissolving 22.61 grams sucrose into water given 969.4 mL of solution. 


1
Expert's answer
2022-03-10T00:10:02-0500

The molar mass of sucrose, C12H22O11=342

Moles of sucrose= 22.61÷342=0.06611

Moles in 1000ml= (0.06611÷969.4)×1000

=0.0682M